OpenStudy (anonymous):

The sum of the measures of two exterior angles of a triangle is 260. What is the measure of the third exterior angle ? A) 80 B)100 C) 110 D) 90

OpenStudy (ranga):

The sum of the exterior angles of any polygon is 360 degrees. A triangle has 3 exterior angles. Two of them add up to 260 degrees. So the third must be ?

OpenStudy (anonymous):

100 lol

OpenStudy (ranga):

yay!

OpenStudy (anonymous):

180-A + 180-B = 260

OpenStudy (anonymous):

Alright you got it lol

OpenStudy (anonymous):

Do you guys want to help me with a couple more problems ? i really need the help i don't understand Geometry at all lol

OpenStudy (ranga):

post them.

OpenStudy (anonymous):

how many sides does a regular polygon have if each exterior angle measures 40 ? A) 9 sides B) 12 sides C) 8 sides D) 11 sides

OpenStudy (ranga):

Same statement made above is true here also. The sum of the exterior angles of ANY polygon is 360 degrees. Let this polygon have N sides. That means it has N exterior angles. N exterior angle adds up to 360. One exterior angle = 360 / N = 40 N = 360 / 40 = 9 sides.

OpenStudy (anonymous):

Find the values of the variables in the parallelogram. The diagram is not drawn to scale . A) x=52, y=60, z=68 B) x=60, y=68, z=52 C) x=60, y=52, z=68 D) x=52, y=68, z=68

OpenStudy (ranga):

The opposite angles of a parallelogram are equal. So z = 68. 52 and y are alternate interior angles because the left and right sides are parallel. So y = 52. The sum of angles of a triangle is 180. So x+y+68 = 180 x+52+68 = 180. x + 120 = 180. x = 60 x=60, y=52, z=68

OpenStudy (anonymous):

For the parallelogram, if m<2=3x-28 and m<4=2x-12, find m<3 . The diagram is not drawn to scale. A) 20 B) 16 C) 160 D)170

OpenStudy (ranga):

Opposite angles of a parallelogram are equal. son angle 2 = angle 4 3x-28 = 2x-12 can you solve for x? Then we can find angle 3.

OpenStudy (anonymous):

is the answer for this problem 16 ? that's what I got but idk if I did it right . lol

OpenStudy (ranga):

No that is the value of x and 16 is correct. Put it in the angle for 4: 2x-12 = 2(16) - 12 = 32 - 12 = 20. In a parallelogram, the angle 4 and angle 3 add to 180. so angle 3 = 180 - 20 = 160.

OpenStudy (anonymous):

ABCD is a parallelogram. The diagram is not drawn to scale. If m<CDA=128, then m<ABC=_____ A) 125 B) 128 C) 52 D) 120

OpenStudy (ranga):

Opposite angles of a parallelogram are equal. Therefore, angle ABC = ?

OpenStudy (anonymous):

128

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

ABCD is a parallelogram. Find AE, if BD= 12, and AC=18 A) 12 B) 18 C) 9 D) 6